wrote a short memo regarding the software reimbursement.
helped Tata with his wireless connection on his notebook.
mn_fit author replied with possible solution. After few tries, able to compiled the program successfully. document all the steps in twiki.
Lee/Sun dropped by with the new system board. System is currently up and running.
troubleshoot Sandip's firefox proble...unable to start it up. Problem traced to a corrupt profile. started firefox with the profile manager and deleted his one but backup his bookmarks first. Use the bookmarks manager to fixed a lot of bad bookmarks.
return the matlab DVD.
11/11:
veteran's holiday
work on Gorham's request regarding su problem to anita account.
11/12:
modified http conf setting and check on imb for access on files and directories.
hep01 crashed again; resurrect the console server; have hep01 console hook up to the console server
troubleshoot Elias HP. Doesn't come on. Have to reseated the memory.
put some web materials for imb account.
replace uhhepq UPS battery; keep the old battery for now because of no shipping label.
have a meeting with Fah and the bes group.
get a quote for UPS battery
 want future /home to be mirrored (raid1)
want future disks to bunched (lvm)
get a quote for esata rack, drives, controller.
11/13:
hep01 crashed again. nothing on the console setenv diag-level to max and turn key to diagnostic. reboot system and reveal some memory errors.
emailed console logs to sun Clayton
updated Margie's adobe program
Lee/Sun called and I told him about the problem. Also emailed the log
Sandip asked about pdf to ppt. Josie told him to install the new Adobe. He will bring his notebook for installation.
research into esata and got a quote from burlystorage.com
11/14:
helped tata with wireless. connection to UH.
helped mwp with wireless. He is not able to connect to UH wireless with his notebook.
Josie is having problem printing to the plotter. Something is wrong with the printer. wait for James to troubleshoot it after spending one hour on it.
un-jammed the hp4100 printer
reinstall adobe acrobat 9.0 on pakvasa's notebook.
Lee/sun replace two memory modules on hep01. will monitor server for now.
phys lab TA complain about pc's slow in psb112. will check it out on Monday.
11/16:
delete tuan's firefox cache - will increase his quota to the next level.
try patch the boss 6.4.0 software - problem with connection to the other side.
hep01 crashed again...emailed console messages to Lee.